OpenStudy (anonymous):

the work done on a current carrying wire by magnetic field is zero...?? there is displacement due to the magnetic force and then too the work done is zero .. Why???

OpenStudy (anonymous):

Magnetic field is nt creating displacement. Displacement is due to its own velocity. Magnetic field just turns the direction.

OpenStudy (anonymous):

its reducing the velocity of wire so work should done .... and it should have been negative

OpenStudy (ghazi):

see the velocity in the wire is developed due to mutual disagreement of magnetic fields by this i mean..when a conductor is placed in a changing magnetic field there is an induced emf in it and hence a current flows (if circuit is complete) so that induced current produces it's own magnetic field ..now due to the difference in magnetic field of conductor and external magnetic field ..the net force is exerted and conductor moves... Now, Emf =B*L*V and i = emf/R= BLV/R and F= i*L*B*sin theta and power= F*Velocity ...so work is done but that is stored in the form of potential energy

OpenStudy (anonymous):

who said that there is no work done ?

OpenStudy (anonymous):

actually work done is zero ... i got it the velocity of electrons have 2 perpendicular components ...(1 due to motion of wire the other due to electric current) the net force is perpendicular to the resultant... and the work done is zero,,,,,